One Number is four more than another. Their sum is sixteen , find both numbers.
Smaller Value Larger Value

Answer by checkley79(3341)

LET X BE THE SMALLER NUMBER
LET X+4BE THE LARGER NUMBER
X+X=4=16
2X=16-4
2X=12
X=6 ANS. FOR THE SMALLER NUMBER
6+4=10 ANS. FOR THE LARGER NUMBER
PROOF:
6+10=16
16=16